a theater has 650 seats. after one sell-out performance, the theater made 11,600. if orchestra seats cost 20 each and balcony seats cost 12, how many orchestra and balcony seats are there in a theater?
Answer by robertb(4012)

let x=# orchestra seats, so 650 - x =# balcony seats.
Then 20x + 12(650 - x) = 11600.
20x - 12x = 11600 - 7800,
8x = 3800,
x = 475 orchestra tickets, and 175 balcony tickets.